OpenStudy (anonymous):

PLEASE HELP!!!! I GIVE MEDALS!!!! The width of a rectangle is 12 units. Can the perimeter x of the rectangle can be 60 units when its length y is 18 units? No, the rectangle cannot have x = 60 and y = 18 because x + y is less than 60. No, the rectangle cannot have x = 60 and y = 18 because x + y is less than 24. Yes, the rectangle can have x = 60 and y = 18 because x = 12 + 2y. Yes, the rectangle can have x = 60 and y = 18 because x = 24 + 2y.

OpenStudy (anonymous):

for a rectangle, perimeter = 2*(length+width) width (x) = 12 inches, if perimeter = 60 60 = 2*(y+12) 60 =2y+24 2y= 60-24 2y =36 y =18inches

OpenStudy (anonymous):

So the answer is which yes?

OpenStudy (anonymous):

The answer would be D